40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 16:16:28 2016, Chris, Omnistructure, VAC, vacuum interlock bypassed 
    Reply  Tue Sep 13 17:47:04 2016, Steve, Update, VAC, pumpdown 79 completed 
       Reply  Wed Sep 14 09:44:09 2016, Steve, Update, VAC, pumpdown 79 completed pd79.pngpd79d2.jpgCCcg.jpg
          Reply  Thu Sep 15 10:36:32 2016, Steve, Update, VAC, pumpdown 79 day 3 vent78for73d.png
Message ID: 12496     Entry time: Thu Sep 15 10:36:32 2016     In reply to: 12492
Author: Steve 
Type: Update 
Category: VAC 
Subject: pumpdown 79 day 3 

IFO pressure 3.7e-5 Torr at new cold cathode InstruTech - Hornet

New items in vacuum:

1, ETMX sus tower with new baked sus wire, EP30-2 epoxied magnets, same at different locations also........ ..........  and 2 ruby wirestandoffs.

2, First Contact cleaned arm test masses only. This technic was a 1st time use in our vacuum system.

3, 50 mm ID green glass baffles at the ends

4, witness mirrors at ETMX and ITMY (old oplev mirrors) We observed a very dusty system: sides of optics, towers and tables were wiped. Hepa tents used at Y arm and BS

5, new pirani, cc gauge and 1.5" right angle valve



The pumpdown started at 9-12-2016

The IFO is at 5e-5 Torr vacuum normal after 73 day at atm.

ITMX needed to be freed and ETMY-UL is still misbehaving occasionally.

New pirani   and cold cathode gauges added at this vent. They were baked at 100 C for 6 hrs under vacuum.

Go to the Vac Rack to read IFO pressure from the gauge itself when Vac. Monitor is blank as it is now !



P1 IFO pressure is 1 mTorr, valve configuration: vacuum normal, annulosses are pumped, RGA is not pumped.

THANKS to Chris !

The shutters can be opened with high power.

No communication error message still exist.

I will reboot c1vac1 and c1vac2 to get gauge communication with medm tomorrow.





Attachment 1: vent78for73d.png  51 kB  | Hide | Hide all
ELOG V3.1.3-